# EventHallusion: Diagnosing Event Hallucinations in Video LLMs

### v1

@@ Introduction
where $S(\cdot)$ represents sample operation on videos, $y_{t}$ is the token to be generated at time step $t$, and $y_{<t}$ denotes the generated token sequence before time step $t$. $y_{t}$ is based on $x$, $V$, and all previously generated tokens $y_{<t}$ by selecting the token with the highest probability in the softmax processed logit distribution.
To construct the temporal modeling distorted video for comparison, we use the sample operation to extract fewer frames, the original sampling operation can be formulated as follows:
$S(V,m)=\left\{I_{t_{1}},I_{t_{2}},\ldots,I_{t_{m}}\right\},$
$t_{i}=\left\lfloor\frac{(i-1)\cdot n}{m-1}\right\rfloor+1,\quad\text{for }i=1,2,\ldots,m$ (2)
where $n$ and $m$ represent the whole frames and sampled frames numbers, we perform the sample operation uniformly along the time. We sample fewer $m^{\prime}<m$ frames $S(V,m^{\prime})$ as the comparison for the contrastive decoding. The whole decoding process can be defined as:
$p_{tcd}$ $\left(y\mid S(V,m),R(S(V,m^{\prime})),x;\theta\right)=$
$\operatorname{softmax}\Big{[}(1+\alpha)\operatorname{logit}_{\theta}\left(y\mid S(V,m),x;\theta\right)$
$\quad-\alpha\operatorname{logit}_{\theta}\left(y\mid S(V,m^{\prime}),x;\theta\right)\Big{]}$ (3)
We adopt the hyperparameter $\alpha$ here to control the strength of contrastive decoding as utilized in (Leng et al. 2024).
As has been considered in (Li et al. 2022), the contrastive decoding process is performed for all tokens to be generated, which will impact the generation of non-visual-related content. Thus, we adopt the adaptive plausible constraint to control tokens for contrastive decoding:
$\begin{array}[]{l}\mathcal{V}_{\text{head }}\left(y_{<t}\right)=\left\{y_{t}\in\mathcal{V}:\right.\\ \left.p\left(y_{t}\mid S(V),x,y_{<t};\theta\right)\geq\beta\max_{w}p\left(w\mid S(V),x,y_{<t};\theta\right)\right\}\end{array}$ (4)
Our TCD is only applied to tokens in the subset $\mathcal{V}_{\text{head }}$, where the hyperparameter $\beta$ in $[0,1]$ controls the truncation of logit distribution.
@@
Models Success Entire Interleave Misleading Overall
Binary Desc. Binary Desc. Binary Desc. Binary Desc.
@@
VideoChatGPT (Maaz et al. 2023) 100.00 14.91 5.50 79.80 3.63 21.57 34.74 47.19 11.59
VideoChat2 (Li et al. 2024b) 83.13 16.67 4.59 58.55 1.55 22.55 24.21 37.90 7.81
PLLaVA (Xu et al. 2024) 99.51 45.61 16.51 68.91 3.11 81.37 51.58 65.53 18.39
LLaMA-VID (Li, Wang, and Jia 2023) 100.00 30.70 16.51 98.96 7.77 43.14 43.16 66.01 18.64
LLaVA-NeXT-Video (Li et al. 2024a) 99.80 49.12 9.17 81.35 4.66 71.57 46.32 69.93 15.87
VILA (Lin et al. 2024) 100.00 53.51 20.18 62.18 17.62 83.33 80.00 65.04 33.25
ShareGPT4Video (Chen et al. 2024) 89.49 11.40 0.00 93.78 5.18 6.86 30.53 49.14 9.82
Video-LLaVA (Lin et al. 2023) 100.00 30.70 8.26 67.36 7.25 41.18 46.32 50.61 16.88
@@
Model Decoding Entire Interleave Misleading Overall
LLaVA-NeXT-Video (Li et al. 2024a) Default 49.12 81.35 71.57 69.93
TCD 50.88 84.46 74.51 72.62
VideoChat2 (Li et al. 2024b) Default 16.67 58.55 22.55 37.90
TCD 17.54 60.62 25.49 39.85
Vila (Lin et al. 2024) Default 53.51 62.18 83.33 65.04
TCD 50.88 70.98 84.31 68.70
Experiments
In this section, we first evaluate multiple prevalent VideoLLMs on our EventHallusion, incorporating both open-sourced and close-sourced MLLMs. Subsequently, we assess our TCD method on our EventHallusion with three VideoLLMs and demonstrate its effectiveness in improving the robustness of VideoLLMs against event hallucinations.
Evaluation on EventHallusion
Evaluated VideoLLMs.
We test 8 open-sourced Video and 2 close-sourced VideoLLMs on our EventHallusion. Open-sourced VideoLLMs comprise VideoChatGPT (Maaz et al. 2023), VideoChat2 (Li et al. 2024b), PLLaVA (Xu et al. 2024), LLaMA-VID (Li, Wang, and Jia 2023), VideoLaVIT (Jin et al. 2024), LLaVA-NeXT-Video (Li et al. 2024a), VILA (Lin et al. 2024), ShareGPT4Video (Chen et al. 2024), and Video-LLaVA (Lin et al. 2023). For close-sourced VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V (Achiam et al. 2023), GPT-4o and Gemini-1.5 (Team et al. 2023), to conduct the systematic comparison of the robustness between open-sourced and close-sourced MLLMs.
@@
Main Results
We exhibit the performance of each VideoLLM in Tab 1 on the binary classification and detailed description matching task. As results have shown, open-sourced VideoLLMs manifest a universal weakness in event hallucinations. In contrast, close-sourced VideoLLMs are more robust against event hallucinations across all categories. We delve into a deeper analysis as follows.
For the evaluation of the sensitivity of VideoLLMs towards video content, the performance on the detailed description matching of Interleave is much lower than the result on the binary classification task, which suggests that VideoLLMs may follow the question as the cue to search rare events. However, they are prone to overlook rare events in their detailed descriptions when no cues are provided. Meanwhile, VideoLLMs have difficulty in processing both binary classification and detailed description matching on Interleave. This indicates the over-reliance of VideoLLMs on language priors. Notably, VideoLLMs show worse capacity on the description matching of Interleave than Entire, which proves that common events can strongly induce VideoLLMs to generate hallucinations than with only rare events.In Misleading, despite their high accuracy on description matching, results on binary classification indicate that VideoLLMs can be affected by questions and generate event hallucinations.
Figure 7: Accuracy of TCD on binary classification of EventHallusion with different numbers of sampled frames in the distorted video. Base indicates the results of standard decoding with 32 frames sampled.
Evaluating TCD on EventHallusion
VideoLLM Baselines.
@@
To evaluate the effectiveness of our TCD in addressing event hallucinations, we adopt LLaVA-NeXT-Video (Li et al. 2024a), VideoChat2 (Li et al. 2024b), and Vila (Lin et al. 2024) as the foundation backbones. LLaVA-NeXT-Video extends the ability of LLaVA (Liu et al. 2024a) based on the projection layer design, while VideoChat2 adopts the Q-Former (Li et al. 2023a) to bridge the gap between videos and LLM. Vila, on the other hand, employs enhanced visual-language pre-training for better video comprehension. Based on these MLLMs, we explore the performance of TCD across different VideoLLM architectures and designs on EventHallusion.
Implementation Details.
@@
We adopt different numbers of extracted frames to construct the original and distorted visual input: 32 and 8 frames for LLaVA-NeXT-Video, 12 and 8 frames for Vila, and 16 and 4 frames for VideoChat2. We follow (Leng et al. 2024) and set $\alpha=1$, $\beta=0.1$. We use greedy search as the basic decoding paradigm.
Main Results
@@
We exhibit the results of the binary classification task on EventHallusion in Tab. 2. As shown in Tab. 2, our TCD demonstrates significant performance across all three VideoLLMs, with increases in overall accuracy, confirming the effectiveness of TCD in mitigating event hallucinations on VideoLLMs with different architectures and visual-language pertaining paradigms. However, while TCD shows consistent and notable improvements on Interleave and Misleading across all three models, the improvement on Entire is minimal, with even a decrease situation observed. This result suggests that entire rare events are a more challenging and hallucination-prone scenario for VideoLLMs.
Effect of Sampled Frames in Distorted Video.
@@
We further explore the effect of sampled frames in distorted video on our TCD. Specifically, we adopt LLaVA-NeXT-Video as the backbone model with 32 frames sampled for the original video input and adjust the numbers of sampled frames in the distorted video. We conduct this experiment on the binary classification task of EventHallusion. As results have shown in Figure 7, TCD maintains a steady improvement in the range of 4 to 16 sampled frames. However, when the number of sampled frames in distorted video decreases from 4 to 2, the performance of TCD deteriorates. We attribute this phenomenon to the excessive destruction of the temporal modeling in the VideoLLM with too few frames incorporated. This leads to drastic changes in the logit of distorted video, thereby affecting the contrastive decoding performance.
Conclusion
@@
In this paper, we systematically investigated the event-related hallucination problem of existing VideoLLMs. To provide a reliable evaluation benchmark to inspect such hallucination degree of VideoLLMs, we curated EventHallusion, which consists of multi-domain videos and annotated questions, both of which are specifically designed to trigger the priors encapsulated in the VideoLLMs themselves. In addition, we also proposed a simple yet effective method, called Temporal Contrastive Decoding (TCD), to alleviate the event-related hallucination problems of existing VideoLLMs in a lightweight plug-and-play manner. We hope our research can benefit further research in the field of reducing the hallucinations of VideoLLMs.
Appendix A A. Automatic Evaluation For Detailed Description Matching
@@
We apply GPT-4o to evaluate the detailed description matching task of our Entire, Interleave, and Misleading categories respectively. For Entire, we prompt GPT-4o to judge if VideoLLMs’ detailed description of the video is consistent with the ground-truth rare event. And for Interleave, we prompt GPT4o to focus on the occurrence of the interleaved rare event. Finally, for Misleading, we prompt GPT4o to assess the consistency of the ground-truth event and search for any unrelated events. In addition, we set up the default system prompt to better guide GPT-4o’s evaluation, which is shown in Figure 8. The detailed prompts are shown in Figure  9, 10, 11, respectively.
@@
Appendix B B. Qualitative Examples
Data examples of EventHallusion.
We present examples of Entire, Interleave, and Misleading in Figure 12, 13, 14, incorporating responses of VideoLLMs. As shown in these examples, VideoLLMs are prone to generate hallucinated answers under all of our settings, indicating their deficiency in addressing event hallucinations.
Examples of GPT-4o’s Evaluation.
We present examples of GPT-4o’s evaluation of detailed description matching task on Entire, Interleave, and Misleading respectively in Figure 15, 16, 17. As shown in these examples, GPT-4o have the capacity to judge the accuracy of video events.
Examples of TCD.
@@
Figure 10: Prompt used for detailed description matching in Interleave.
@@
Figure 13: Examples of Interleave with responses of VideoLLMs.
@@
Figure 16: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Interleave.

### v2

@@ 4 Temporal Contrastive Decoding
where $y_{t}$ is the token to be generated at time step $t$, and $y_{<t}$ denotes the generated token sequence before time step $t$. $y_{t}$ is based on $x$, $V$, and all previously generated tokens $y_{<t}$ by selecting the token with the highest probability.
When constructing a contrastive counterpart, we chronologically downsample the original video as shown in the Fig.5, aiming to weaken the temporal cues while emphasizing object cues. Therefore, the contrastive counterpart’s probability prediction $z_{t}^{con}$ at timestep $t$ can be formulated as:
$z_{t}^{con}=\operatorname{logit}_{\theta}\left(y_{t}\mid S(V),x,y_{<t}\right)$ (2)
where $S(\cdot)$ represents the temporal downsampling operation. With the generated $z_{t}^{ori}$ and $z_{t}^{con}$, the modulated probability distribution $p_{t}$ at the current timestep can be calculated as:
$p_{t}=\operatorname{softmax}\Big{[}(1+\alpha)z_{t}^{ori}-\alpha z_{t}^{con}\Big{]}$ (3)
where $\alpha$ is the hyperparamter to control the contrastive strength during the decoding. However, applying such a contrastive process throughout the entire decoding stage would lead to significant model degradation, as it is not always consistent for $z_{t}^{ori}$ and $z_{t}^{con}$ to have mismatched distribution. When $z_{t}^{ori}$ and $z_{t}^{con}$ are consistent and correct, employing Eq.3 to modulate the probability impairs the model’s judgment. Therefore, we introduce another hyperparameter $\beta$ to constrain the token set used for TCD. The motivation behind this design is twofold: when the model exhibits high confidence in a specific token, we conduct TCD on a smaller subset of tokens with similarly high confidence. Conversely, when the model’s confidence is distributed more evenly across tokens, we apply a lower threshold, enabling TCD on a broader range of tokens.
@@
Models Entire Mix Misleading Overall
Binary Desc. Binary Desc. Binary Binary Desc.
@@
VideoChatGPT [32] 14.91 5.50 56.99 3.63 21.57 36.43 4.30
VideoChat2 [21] 16.67 4.59 12.44 1.55 22.55 16.14 2.64
PLLaVA [39] 45.61 16.51 58.55 3.11 81.37 60.64 6.05
LLaMA-VID [23] 30.70 16.51 73.58 7.77 43.14 54.03 10.90
LLaVA-NeXT-Video [19] 49.12 9.17 70.47 4.66 71.57 64.79 6.29
VILA [26] 53.51 20.18 58.55 17.62 83.33 63.33 18.54
ShareGPT4Video [8] 11.40 0.00 67.88 5.18 6.86 49.14 9.82
Video-LLaVA [25] 30.70 8.26 57.51 7.25 41.18 45.97 7.62
@@
Model Decoding Entire Mix Misleading Overall
LLaVA-NeXT-Video [19] Default 49.12 70.47 71.57 64.79
TCD 50.88 73.58 74.51 67.48
VideoChat2 [21] Default 16.67 47.67 22.55 32.76
TCD 17.54 49.74 25.49 34.72
VILA [26] Default 53.51 58.55 83.33 63.33
TCD 50.88 66.32 84.31 66.50
5 Experiments
In this section, we first evaluate multiple prevalent VideoLLMs on our EventHallusion, including both open-sourced and private MLLMs. Subsequently, we examine the effectiveness of our proposed TCD, and also discuss its more specific design choices.
5.1 Evaluation on EventHallusion
5.1.1 Selected VideoLLMs for Evaluation
We test 8 open-sourced Video and 2 close-sourced VideoLLMs on our EventHallusion. Open-sourced VideoLLMs comprise VideoChatGPT [32], VideoChat2 [21], PLLaVA [39], LLaMA-VID [23], VideoLaVIT [16], LLaVA-NeXT-Video [19], VILA [26], ShareGPT4Video [8], and Video-LLaVA [25]. For close-sourced VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V [1], GPT-4o and Gemini-1.5 [35], to conduct the systematic comparison of the robustness between open-sourced and close-sourced MLLMs.
@@ 5.1.1 Selected VideoLLMs for Evaluation
5.1.2 Performance Analysis
We exhibit the performance of each VideoLLM in Tab 1 on the binary classification and detailed description matching task. As the results indicate, open-source VideoLLMs exhibit a consistent weakness in event hallucinations. In contrast, closed-source VideoLLMs demonstrate greater robustness against event hallucinations across all categories. A deeper analysis is provided below: (1) In evaluating the susceptibility of VideoLLMs to vision-language correlation biases (i.e, “Entire” and “Mix” in Tab.1), most VideoLLMs perform worse when confronted with videos containing a mix of common and rare events, compared to those containing only rare events. This result indicates that frequent events can distract the model’s attention, imposing higher challenges to the model. (2) In evaluating the susceptibility of VideoLLMs to language priors (i.e., “Misleading” in Tab.1), closed-source models demonstrate overwhelming advantages compared to open-source models, suggesting that these closed-source VideoLLMs are built upon language foundation models that are significantly more powerful than their open-source counterparts.
5.2 TCD Evaluation
5.2.1 VideoLLM Baselines.
@@ 5.2 TCD Evaluation
To evaluate the effectiveness of our TCD in addressing event hallucinations, we adopt LLaVA-NeXT-Video [19], VideoChat2 [21], and VILA [26] as the foundation backbones. LLaVA-NeXT-Video extends the ability of LLaVA [28] based on the projection layer design, while VideoChat2 adopts the Q-Former [20] to bridge the gap between videos and LLM. VILA, on the other hand, employs enhanced visual-language pre-training for better video comprehension. Based on these MLLMs, we explore the performance of TCD across different VideoLLM architectures and designs on EventHallusion.
5.2.2 Implementation Details.
@@
We adopt different numbers of extracted frames to construct the original and distorted visual input: 32 and 8 frames for LLaVA-NeXT-Video, 12 and 8 frames for VILA, and 16 and 4 frames for VideoChat2. We use greedy search as the basic decoding paradigm.
5.2.3 Main Results Analysis
@@
We exhibit the results of the binary classification task on EventHallusion in Tab.2. As shown in Tab.2, our TCD demonstrates significant performance across all three VideoLLMs, with increases in overall accuracy, confirming the effectiveness of TCD in mitigating event hallucinations on VideoLLMs with different architectures and visual-language pertaining paradigms. However, while TCD shows consistent and notable improvements on Mix and Misleading across all three models, the improvement on Entire is minimal, with even a decrease situation observed. This result suggests that entire rare events are a more challenging and hallucination-prone scenario for VideoLLMs.
5.2.4 Ablations on Contrastive Sample Constructions.
We also discuss some different alternative strategies when constructing the counterpart of the original video for performing our TCD method as shown in Tab.3. Specifically, we compare our base TCD method with adding Gaussian noises to video frames when employing TCD method (denoted as TCD+noises). As results have shown in Tab.3, While both TCD and TCD+noises demonstrate improvement on regular decoding. However, TCD+noises does not exhibit large improvements on simply using TCD method.
Table 3: Ablation study of different strategies when constructing the contrastive counterpart of the original video. The best results of each category are bolded.
Method Entire Mix Misleading Overall
Baseline 49.12 70.47 71.57 64.79
TCD 50.88 73.58 74.50 67.48
TCD+noise 51.75 73.58 74.50 67.73
Table 4: Ablation study of $\alpha$ in our TCD. The best results of each category are bolded.
$\alpha$ Entire Mix Misleading Overall
0.25 50.00 73.56 73.53 66.99
0.5 50.88 73.56 74.51 67.48
0.75 50.88 71.50 74.51 66.50
1.0 49.12 64.25 74.51 62.59
Table 5: Ablation study of $\beta$ in our TCD. The best results of each category are bolded.
$\beta$ Entire Mix Misleading Overall
0.1 50.88 73.58 74.51 67.48
0.2 50.88 73.58 74.51 67.48
0.5 50.88 74.09 74.51 67.73
0.5 50.00 74.61 74.51 67.73
5.2.5 Ablations on Hyper-parameters of TCD
We conduct ablation studies to investigate the effects of hyper-parameters $\alpha$ and $\beta$ on our TCD method as shown in Tab.4 and Tab.5 For the first hyper-parameter $\alpha$, with the growth of $\alpha$, TCD applies a more fierce contrast between normal and distorted video input. As results are shown in Tab.4, when $\alpha$ is 0.5, TCD exhibits best performance. While for $\beta$, the threshold used to control the token set for TCD does not exhibit a strong impact on the performance of TCD on our EventHallusion Benchmark.
5.2.6 Effect of Sampled Frames in Distorted Video.
@@ 5.2.5 Ablations on Hyper-parameters of TCD
We further explore the effect of sampled frames in distorted video on our TCD. Specifically, we adopt LLaVA-NeXT-Video as the backbone model with 32 frames sampled for the original video input and adjust the numbers of sampled frames in the distorted video. We conduct this experiment on the binary classification task of EventHallusion. As results have shown in Tab.6, TCD maintains a steady improvement in the range of 4 to 16 sampled frames. However, when the number of sampled frames in distorted video decreases from 4 to 2, the performance of TCD deteriorates. We attribute this phenomenon to the excessive destruction of the temporal modeling in the VideoLLM with too few frames incorporated. This leads to drastic changes in the logit of distorted video, thereby affecting the contrastive decoding performance.
Table 6: Ablation study of extracted video frames in the distorted group. The best results of each category are bolded.
Frames Entire Mix Misleading Overall
1 48.26 68.39 74.51 64.30
4 50.00 73.06 76.47 67.48
8 50.88 73.58 74.51 67.48
16 50.88 73.06 74.51 66.75
Figure 6: Qualitative Results of the TCD method on our EventHallusion Benchmark.
5.3 Qualitative Results
We exhibit an example of our question in EventHallusion. As shown in Fig.6, VideoLLM can be affected and generate hallucination answers, while our TCD method can alleviate this problem. Due to the page limit, we provide more examples in our suuplementary materials.
6 Conclusion
@@
In this paper, we systematically investigated the event-related hallucination problem of existing VideoLLMs. We curated EventHallusion, which consists of multi-domain videos and annotated questions, both of which are specifically designed to trigger the priors encapsulated in the VideoLLMs themselves. In addition, we also proposed a simple yet effective method, called Temporal Contrastive Decoding (TCD), to alleviate the event-related hallucination problems.
Supplementary Material
1 Comparison with other video hallcination benmarks.
We exhibit the comparison between HallusionBench, VideoHallucer, and our EventHallusion in Figure 7. As shown in Figure 7, compared to HallusionBench and Videohallucer emphasize assessing the models’ sensitivity to events’ chronological orders, our proposed EventHallusion focuses on analyzing the event-related language prior and vision-language correlation bias encapsulated in existing VideoLLMs.
2 Automatic Evaluation For Detailed Description Matching
@@
We apply GPT-4o to evaluate the detailed description matching task of our Entire, Mix, and Misleading categories respectively. For Entire, we prompt GPT-4o to judge if VideoLLMs’ detailed description of the video is consistent with the ground-truth rare event. And for Mix, we prompt GPT4o to focus on the occurrence of the rare-common mixed event. Finally, for Misleading, we prompt GPT4o to assess the consistency of the ground-truth event and search for any unrelated events. In addition, we set up the default system prompt to better guide GPT-4o’s evaluation, which is shown in Figure 8. The detailed prompts are shown in Figure  9, 10, 11, respectively.
Figure 7: Comparison between HallusionBench, VideoHallucer, and our EventHallusion. We use Green to highlight the content of the original video/question, and red to denote refactored video/question content to assess the model’s event-related hallucinations.
@@
3 Qualitative Examples
3.1 Data examples of EventHallusion.
We present examples of Entire, Mix, and Misleading in Figure 12, 13, 14, incorporating responses of VideoLLMs. As shown in these examples, VideoLLMs are prone to generate hallucinated answers under all of our settings, indicating their deficiency in addressing event hallucinations.
3.2 Examples of GPT-4o’s Evaluation.
We present examples of GPT-4o’s evaluation of detailed description matching task on Entire, Mix, and Misleading respectively in Figure 15, 16, 17. As shown in these examples, GPT-4o have the capacity to judge the accuracy of video events.
3.3 Examples of TCD.
@@ 3 Qualitative Examples
Figure 10: Prompt used for detailed description matching in Mix.
@@
Figure 13: Examples of Mix with responses of VideoLLMs.
@@
Figure 16: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Mix.